OpenStudy (astrophysics):

DE series @ganeshie8

OpenStudy (astrophysics):

If I'm given a differential equation and have to solve it via power series, would I then just use Airy's equation, \[y = \sum_{n=0}^{\infty} a_n(x-1)^n\] as my DE is \[y''+\frac{ 2 }{ (4-x^2) }y=0\]

OpenStudy (astrophysics):

Or wait I guess I can use \[\sum_{n=0}^{\infty} a_n x^n\] damn I'm dumb xD

OpenStudy (astrophysics):

That's like the assumption I should always be making it looks like haha

OpenStudy (astrophysics):

Then \[y = \sum_{n=0}^{\infty} a_n x^n\] \[y' = \sum_{n=1}^{\infty} a_{n+1} nx^{n-1}\] \[y''= \sum_{n=2}^{\infty} a_nn(n-1)x^{n-2}\] ehh ehhh now plug into the equation

OpenStudy (astrophysics):

Oh \[x_0 = 0 \]

OpenStudy (astrophysics):

\[\sum_{n=2}^{\infty} a_n n(n-1)x^{n-2}+\frac{ 2 }{ (4-x^2) }\sum_{n=0}^{\infty} a_nx^n\] what is this

OpenStudy (astrophysics):

Ok shift index now here we go, this is where I'll make a mistake

OpenStudy (astrophysics):

i = n-2 guaranteed I'm calling it then n = i+2

OpenStudy (astrophysics):

dont reply im gonna see this hold onnn

OpenStudy (astrophysics):

noo dont

OpenStudy (astrophysics):

leave

OpenStudy (dan815):

yep just make the indeces match up

OpenStudy (astrophysics):

Oh yeah I think it's looking rather fancy, lets see if you agree

OpenStudy (astrophysics):

\[\sum_{n=0}^{\infty} a_{n+2}(n+2)(n+1)x^n+\frac{ 2 }{ (4-x^2) } \sum_{n=0}^{\infty} a_n x^n\]

OpenStudy (astrophysics):

err

OpenStudy (astrophysics):

I need a recurrence relationship

OpenStudy (astrophysics):

What to do with 2/(4-x^2)

OpenStudy (astrophysics):

oh I guess, \[\sum_{n=0}^{\infty} [a_{n+2}(n+2)(n+1)+a_n \frac{ 2 }{ (4-x^2) }]x^n=0\]

OpenStudy (astrophysics):

omggg you dummy

OpenStudy (astrophysics):

don't need to put it in standard form

OpenStudy (astrophysics):

!@%$!@$!@$

imqwerty (imqwerty):

m totally blank i jst knw how to solve this kinda things- y'-y=0 lol

OpenStudy (astrophysics):

I have to erase some stuff now lol, and it's ok

OpenStudy (astrophysics):

It was making no sense, because there shouldn't be any x's err ok let me see again

ganeshie8 (ganeshie8):

hey still working on this ?

OpenStudy (astrophysics):

Yup, there's a few parts to this problem, I see my mistake so I'm trying to fix that right now

OpenStudy (astrophysics):

\[(4-x^2)\sum_{n=2}^{\infty} a_n n(n-1)x^{n-2}+2 \sum_{n=0}^{\infty} a_n x^n=0\] this is what I have, I'm trying to find a recurrence relationship, so lets see what I get haha

OpenStudy (astrophysics):

Ok I think I have to split the first series in two because that -x^2 will mess things up

OpenStudy (astrophysics):

Oh perfect, I think that does work!

OpenStudy (astrophysics):

\[4\sum_{n=2}^{\infty}a_nn(n-1)x^{n-2}-\sum_{n=2}^{\infty} a_nn(n-1)x^{n-2+2} + 2 \sum_{n=0}^{\infty} a_n x^n=0\] ok lol

OpenStudy (astrophysics):

Now I can shift the index for those two series

OpenStudy (astrophysics):

I hope that looks right

ganeshie8 (ganeshie8):

Looks good to me... you ust need to shift the index to first term

ganeshie8 (ganeshie8):

\[4\sum_{n=2}^{\infty}a_nn(n-1)x^{n-2}-\sum_{n=2}^{\infty} a_nn(n-1)x^{n-2+2} + 2 \sum_{n=0}^{\infty} a_n x^n=0\] is same as \[4\sum_{n=2}^{\infty}a_nn(n-1)x^{n-2}-\sum_{n=0}^{\infty} a_nn(n-1)x^{n-2+2} + 2 \sum_{n=0}^{\infty} a_n x^n=0\]

ganeshie8 (ganeshie8):

Notice that \[ \sum\limits_{n=0}^5 n(n-1) = \sum\limits_{n=2}^5 n(n-1) \]

OpenStudy (astrophysics):

Yeah I see, that's what I sort of noticed from shifting the index, you should actually get the same result for when you change the index as you would with the original series, just hard to tell sometimes without plugging in random numbers haha

OpenStudy (astrophysics):

Well I guess it's not very random...

OpenStudy (astrophysics):

But how did you know right away the first one was correct :O

ganeshie8 (ganeshie8):

idk the recurrence relation looks nasty..

OpenStudy (astrophysics):

lol ok one second let me see what I get

ganeshie8 (ganeshie8):

\[4(n+1)(n+2)a_{n+2}-[n(n+1)-2]a_n=0\]

ganeshie8 (ganeshie8):

do double check...

OpenStudy (dan815):

http://prntscr.com/8yuey8

OpenStudy (dan815):

http://prntscr.com/8yuf60

OpenStudy (astrophysics):

\[4a_{n+2}(n+2)(n+1)-a_nn(n-1)+2a_n=0\]

OpenStudy (dan815):

ya thats right

OpenStudy (astrophysics):

Yeah so -2 because when you factor out the a_n ok ok

OpenStudy (dan815):

ya

OpenStudy (astrophysics):

So this would be the recurrence relation given the point x0

ganeshie8 (ganeshie8):

yes just solve it and you're done ! ;)

OpenStudy (astrophysics):

Hey, so these are the solutions, and we don't have to use like variation of parameter to find the general solution, none of that characteristic equation business eh

ganeshie8 (ganeshie8):

It is a very simple linear equation, you could have solved it by that integration factor thingy... but you wanted to solve it using power series..

OpenStudy (astrophysics):

Yeah haha ok, it just wanted me to find the recurrence relation now 4 terms...blah blah, then find general term in each solution

OpenStudy (astrophysics):

Hey quick question, remember the method of undetermined coeff, the guess is basically like using an integrating factor, I think that's why you can't have duplicate stuff, thought I'd let you know if you didn't already...or you can ignore it if it doesn't make sense

ganeshie8 (ganeshie8):

for lienar equations, integrating factor is a clever way to make use of reverse product rule... what has that anythign to do with the guess for a particular solution in nonhomogeneous eqn ?

ganeshie8 (ganeshie8):

here is the solution generated by wolfram... this time, its really a very smart solution!

OpenStudy (astrophysics):

I mean if you have something like y'+y=2e^-t, which then gives you \[y=2te^{-t}+ce^{-t}\] and this gives you a clue on your guess, just by looking at this solution

OpenStudy (astrophysics):

Oooh

ganeshie8 (ganeshie8):

got you :)

OpenStudy (empty):

I am looking at the problem I might have found an alternate way to an answer but it's quite hairy in its own way lol

OpenStudy (astrophysics):

I don't think I've used integrating factor like this on a second order? That's pretty cool

ganeshie8 (ganeshie8):

hey me neither, i got confused earlier...

OpenStudy (astrophysics):

Looks like it has the same amount of work as series

ganeshie8 (ganeshie8):

yeah

OpenStudy (astrophysics):

So to find the first four terms, I have \[a_{n+2} = \frac{ a_n[n(n-1)-2] }{ 4(n+2)(n+1) }\] lol whyyyy

OpenStudy (empty):

Tell me why this doesn't work cause it seems too simple to do it this way: Start from the DE: \[(x^2-4)y''-2y=0\] substitute \(x^2-4=u^2\) \[u^2y''-2y=0\] Now we see that this is an Euler equation, \(y=u^n\) we solve for n, \[n(n-1)u^n-2u^n=0\] \[n^2-n-2=(n-2)(n+1)=0\] so we have the general solution \[y=c_1u^2+c_2u^{-1}\] substitute in \[u=\sqrt{x^2-4}\] \[y=c_1(x^2-4) + \frac{c_2}{ \sqrt{x^2-4}}\] I guess I should check this now.

OpenStudy (astrophysics):

That's weird, because there's no y' term i think which sort of messes it up

OpenStudy (astrophysics):

So I don't think Euler's formula will be applicable for this one?

OpenStudy (empty):

That doesn't matter, I'll show the steps: \[u^2y''-2y=0\] is the same as \[u^2y''+0uy'-2y=0\] Take: \[y=u^n\]\[y'=nu^{n-1}\]\[y''=n(n-1)u^{n-2}\] Plug it in accordingly: \[u^2(n(n-1)u^{n-2})-2u^n=0\]divide out the \(u^n\) term in common to get a quadratic \[n^2-n-2=(n-2)(n+1)=0\] so we have the general solution which are the roots \[y=c_1u^2+c_2u^{-1}\]

OpenStudy (astrophysics):

wait so is y = u^n like your substitution where you then calculate y' and y'' in terms of dy/dx and d^2y/dx^2 idk what you're doing

OpenStudy (astrophysics):

qwerty quit logging out haha making me lag when you come as anonymous user

OpenStudy (empty):

Well since I made the substitution that x is a function of u now, y is a function of u as wel so I'm taking the derivative with respect to u, \[x(u)\] \[y(x(u))\] Well I guess we should just check that \[ y=c_1(x^2-4) + \frac{c_2}{ \sqrt{x^2-4}}\] is a solution to the differential equation before I talk anymore lol

OpenStudy (empty):

It's strange cause the first part is a solution but the second part doesn't seem to be, I think there's something to do with the square root, perhaps I need to be doing like \(\pm\) in front of it? Interesting.

OpenStudy (empty):

At any rate, once you have seen that the first one is a solution, we can try looking for another solution that's a multiple of the original solution by plugging in: \[y=(x^2-4)f(x)\] this reduces our differential equation (if i didn't mess up) to: \[4xf'+2(x^2-4)f''=0\] Now you might notice after you distribute this looks interesting: \[4xf'+2x^2f''=8f''\] That left hand side looks an awful lot like the product rule, so with a little playing around you can see that it's just going to be : \[(2x^2f')'=4xf'+2x^2f''\] so we plug it in: \[(2x^2f')'=8f''\] Now divide both sides by 2 and let's integrate! \[x^2f' = 4f' +C\] This is looking good. \[f'(x^2-4)=C\] \[f=\int \frac{C dx}{x^2-4}\] Yay the rest from here should be pretty straight forward and it looks like it's gonna match. Whew.

OpenStudy (empty):

Note: Anything that you see here that seems crazy is something I've seen someone else do before. Specifically when I try looking for another solution by simplifying the differential equation with this multiplication, this is a trick and it's sorta related to asymptotic analysis which is just a fancy way of saying you solve the differential equation after making simplifications like as x gets infinite or really small, we might look at: \[y'' + \frac{2}{4-x^2}y=0\] so as x is large, that 4 in the denominator is less important, so we can solve the much easier equation: \[y''-\frac{2}{x^2}y=0\] and then taking the solution to this differential equation, multiplying it by something like we did earlier to try to reduce the complexity of the differential equation. Similarly we could have looked at small x where now the importan term is: \[y''+\frac{1}{2}y=0\] and done it this way as well. So that first one which is for large x ends up getting us an euler equation which is why I thought to use this trick while the second one ends up in a sinusoid that didn't take me anywhere. --- The other trick I used was reversing the product rule weirdly. My partial differential equations prof did this all the time and after a while you get used to it. A pretty awesome trick I like to use is this one: \[y'' = \frac{d y'}{dx} = \frac{dy'}{dy}\frac{dy}{dx} = \frac{dy'}{dy} y'\] which can make things separable.

OpenStudy (astrophysics):

Cool stuff man, thanks for sharing haha, I'll have to go over this again, as I'm sort of multitasking and doing my assignment